OpenStudy (anonymous):

A volcanic plug shaped like a cone has a diameter of 24 km and a height of 3 km. What is the volume of the plug? Use 3.14 to approximate pi, and express your final answer in hundredths. km3

OpenStudy (anonymous):

@amistre64 @dtan5457 @Mehek14 @texaschic101 @Michele_Laino

Mehek (mehek14):

do you know the formula?

OpenStudy (anonymous):

V=3.14r^2

Mehek (mehek14):

that's for a circle

Mehek (mehek14):

I need the formula for a cone

Mehek (mehek14):

\(\dfrac{1}{3}{πr^2}\)

OpenStudy (anonymous):

whats the squares?

Mehek (mehek14):

refresh the page

Mehek (mehek14):

got it?

OpenStudy (anonymous):

yea

Mehek (mehek14):

so you have a diameter of 24 you need the radius so 24/2

Mehek (mehek14):

divide

OpenStudy (texaschic101):

radius is half the diameter

OpenStudy (anonymous):

12

OpenStudy (texaschic101):

V = 1/3 * (pi) * r^2 * h pi = 3.14 r (radius) = 1/2 diameter.....24/2 = 12 h = 3 V = 1/3 * (3.14) * (12^2) * 3 V = 1/3 * 3.14 * 144 * 3 V = 1/3 * 1356.48 V = 1356.48/3 V = 452.16

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

anytime :)